OpenStudy (anonymous):

E=mc2 is one of the most recognizable equations in the world. In this task, you’ll attempt to understand it better. Let’s start with what the letters mean. The letter E stands for energy, the letter m stands for mass, and the letter c stands for the speed of light. The meaning of this equation is that you can completely convert mass into energy. a. Rewrite the equation to solve for the speed of light, c. Use rational exponents instead of roots.

OpenStudy (greatlife44):

\[E = mc^{2}\] we need to solve for c

OpenStudy (greatlife44):

we can do this by first isolating c

OpenStudy (greatlife44):

\[\frac{ E }{ m } = c^{2}\]

OpenStudy (greatlife44):

now after this. we have c^{2} the opposite would be to take the square root, Let's do that. \[\sqrt{ c^{2}} = \sqrt{\frac{ E }{ m }}\] \[\sqrt{ \frac{ E }{ M }} = c\] or \[\frac{ \sqrt{E} }{ \sqrt{m} } = c \]

OpenStudy (greatlife44):

if you notice whatever you do to one side you need to do to the other side.

OpenStudy (anonymous):

cada uno es fresco